What will be the de broglie wavwlength when the kinetic energy of body increases by 5 times



1
Expert's answer
2021-09-21T11:20:06-0400

Gives

E'=E

"\lambda'=?"

"\lambda=\frac{h}{\sqrt{2mE}}"

"\frac{\lambda'}{\lambda}=\frac{\sqrt{2mE}}{\sqrt{2mE'}}"

"\frac{\lambda'}{\lambda}=\frac{\sqrt{E}}{\sqrt{E'}}"

"\frac{\lambda'}{\lambda}=\frac{\sqrt{E}}{\sqrt{5E}}"

"\frac{\lambda'}{\lambda}=\frac{\sqrt{1}}{\sqrt{5}}"

"{\lambda'}=\frac{{\lambda}}{\sqrt{5}}"